OpenStudy (venomous.penguin):

A trapezoid has base lengths 4 in. and 8 in. and height 8 in. What happens to the area and the perimeter of the trapezoid when its bases and height are all multiplied by 1/8 (one-eighth)? The perimeter is multiplied by 1/4. The area is multiplied by 1/8. The perimeter is multiplied by 1/4. The area is multiplied by 1/16. The perimeter is multiplied by 1/64. The area is multiplied by 1/64. The perimeter is multiplied by 1/8. The area is multiplied by 1/64.

OpenStudy (tubaairfan):

lol

OpenStudy (venomous.penguin):

Are you able to help me?

OpenStudy (mww):

This is a question about similarity. Similar figures are simply scaled up or down copies of the original figure. Here the new figure has a scale factor of 1/8 (since each new side is exactly 1/8 of the old) Therefore the perimeter must also be 1/8 of the original. The area though would be (1/8)^2 of the original = 1/64 because area is two dimensions, so multiplying 1/8 by 1/8 gives 1/64 for the two dimensions used. In general if a figure is a scale factor of a:b of the original: Then each side (and the perimeter) is a factor a:b of the original sides The area scale factor is the square of this so a^2:b^2 \[Area_{new} = (a^2:b^2) Area_{old}\] Volume would be the cube of this so a^3:b^3. You can also test this by seeing your actual sides Here you have

OpenStudy (zzr0ck3r):

can you explain what \((a^2:b^2)\) is

OpenStudy (venomous.penguin):

Yes what is that formula? and the answer would be -> The perimeter is multiplied by 1/8. The area is multiplied by 1/64

OpenStudy (mww):

a:b is the scale factor between the new and old figures. so in this case it is 1:8 (a:b)^2 is then the square of this which is 1:64

OpenStudy (mww):

As an example: